# An Empirical Evaluation of Pre-trained Large Language Models for Repairing Declarative Formal Specifications

### v1

@@ 5.4. Results for RQ4: Characteristics of Failed Repairs
Figure 4 illustrates the distribution of failed iterations across distinct settings and provides a comprehensive breakdown of failure reasons. Predominantly, failures arise when the proposed repairs fail to fulfill assertions, resulting in the generation of counterexamples. Additionally, syntax errors in proposed repairs impede the compilation process by Alloy Analyzer. Settings 1-3, employing GPT-4-32k, manifest instances of failures attributed to incorrect message formats, a phenomenon absent in Settings 4-6 utilizing GPT-4-Turbo. Such disparities suggest discrepancies in the compliance of LLM responses with the JSON format mandated by the alloy_analyzer_tool. Categories such as ”Repetition” and ”No instances” exhibit lower prevalence. ”Repetition” denotes instances where proposed repairs mimic defective models, while ”No instances” signify scenarios where proposed models fail to generate instances, indicative of an incapacity to meet model constraints.
Figure 5. Repair progress in Setting-6 (GPT-4-Turbo and Auto-Feedback) across iterations. Percentage numbers show transition volumes between statuses in subsequent iterations. Cumulative counts of fixed models follow each ’Fixed’ block.
@@
Figure 5 tracks outcomes across all iterations in Setting-6, revealing the evolving nature of failure categories during the iterative prompting process. While instances of ”Counterexample” failures are effectively addressed, challenges persist in rectifying issues classified under ”Repetition.” Moreover, a discernible transition of cases between distinct failure categories underscores the dynamic nature of failure characteristics throughout the repair endeavor.
@@
Analysis of failure characteristics during the repair process highlights common issues such as proposed repairs failing assertions and syntax errors, with observed transitions between failure categories over iterative prompting.
5.5. Results for RQ5: Repair Costs
@@
To assess the financial implications of utilizing the APR pipeline with different LLMs, we consider two primary cost factors: (i) time taken for bug resolution and (ii) monetary expenses associated with token consumption, based on OpenAI’s pricing as of March 2024.
@@
Time Analysis: We measure the time taken for bug resolution using the results presented in Table 6, which leverages GPT-4 models. Figure 6 illustrates the runtime analysis, contrasting fixed and unfixed scenarios. Across different settings, the median duration required for bug repair varies from 37.3 seconds in Setting-1 to 120.3 seconds in Setting-3. Notably, the maximum time spent on repairs reached 493.28 seconds in Setting-5. Conversely, the median duration associated with unfixed bugs amounted to 273.8 seconds and 392.2 seconds in Settings-1 and 4, respectively.
@@
Figure 6. Distribution of running time across all settings for fixed bugs (left) and unfixed bugs (right).
@@
Monetary Costs: Figure 7 presents the monetary costs incurred under each setting for both fixed and unfixed models. Settings utilizing GPT-4-Turbo show the lowest costs for both fixed and unfixed models, attributed to the cheaper cost per token for GPT-4-Turbo compared to GPT-4-32k. Additionally, settings employing the prompt-agent (Settings-3 and 6) exhibit marginally higher costs than their counterparts utilizing the same LLM.
@@
Figure 7. Monetary distribution across all settings for fixed bugs (left) and unfixed bugs (right).
The analysis of repair costs in RQ5 reveals variations in both time and monetary expenses across different settings, with GPT-4-Turbo exhibiting the lowest costs for bug resolution, and settings utilizing the prompt-agent showing slightly higher expenses.
6. Threats to Validity
@@
External Validity. The generalizability of our findings may be limited by the specific characteristics of the benchmarks used. To address this, we utilized diverse benchmark suites and evaluated the performance of the repair pipeline across various scenarios. Furthermore, we documented the experimental setup comprehensively to facilitate replication and external validation of our results.
@@
7. Related Work
@@
Several recent studies have explored the integration of Large Language Models (LLMs) into software engineering tasks, particularly in the realm of program repair. Xia et al. (Xia and Zhang, 2023) enhanced automated program repair (APR) performance by incorporating test feedback into prompts, while Kang et al. (Kang et al., 2023) enabled LLMs to utilize a debugger for information gathering and patch generation. Additionally, Fuzz4All (Xia et al., 2024) leverages LLMs as an input generation and mutation engine, employing an auto-prompting phase to generate concise input prompts. RepairAgent (Bouzenia et al., 2024) employs LLMs, agents, and dynamic prompts for APR, albeit in the context of repairing Java applications. TestPilot (Schäfer et al., 2024) utilizes LLMs for generating unit test cases for JavaScript, employing a few-shot learning approach to refine prompts with failing tests and error messages.
@@
In contrast to these efforts, our study focuses specifically on using LLMs for automated specification repair (ASR), a less explored area in software engineering. We adopt an approach similar to Fuzz4All’s auto-prompting phase, albeit with modifications to suit the requirements of repairing Alloy declarative specifications. Moreover, while existing works predominantly use LLMs for program repair in imperative languages like Java and JavaScript, our study extends the application of LLMs to the domain of Alloy specifications, addressing a gap in the literature regarding ASR techniques for declarative languages.
8. Conclusion
@@
In this study, we explore the potential of pre-trained LLMs to facilitate the repair of Alloy specifications, taking into account recent advancements in LLMs, including the use of agents, feedback mechanisms, zero-shot learning capabilities, and auto-prompting techniques. The comparative evaluation highlights the superior performance of the GPT-4 model family over GPT-3.5-Turbo, underscoring their promising applicability. The investigation reveals that employing a multi-agent repair pipeline enhances the repair process, albeit with a marginal increase in token consumption. Overall, the findings of this research indicate a positive outlook for the application of LLMs in automated program repair for declarative specifications, particularly through repair pipelines that integrate contemporary innovations in the field of LLMs.

### v2

@@ 5.3 Results for RQ3: Characteristics of Failed Repairs
This research question explores the underlying causes of unsuccessful repair iterations. Figure 4 illustrates the distribution of failed iterations in different settings and provides a comprehensive breakdown of the reasons for the failure. Predominantly, failures arise when the proposed repairs fail to fulfill the assertions, resulting in the generation of counterexamples. Additionally, syntax errors in proposed repairs impede the compilation process by Alloy Analyzer. Settings 1-3, employing GPT-4-32k, manifest instances of failures attributed to incorrect message formats, a phenomenon absent in Settings 4-6 utilizing GPT-4-Turbo. Such disparities suggest discrepancies in the compliance of LLM responses with the JSON format mandated by the alloy_analyzer_tool. Categories such as “Repetition” and “No instances” exhibit lower prevalence. “Repetition” denotes instances where proposed repairs same as the supplied defective models, while “No instances” signify scenarios where proposed models fail to generate instances, indicative of an incapacity to meet model constraints.
Figure 4: Error types observed in failed repair iterations across all settings
@@
Figure 5 tracks the results across all iterations in Setting-6, revealing the evolving nature of failure categories during the iterative prompting process. Although instances of “Counterexample” failures are effectively addressed, challenges persist to rectify issues classified under “Repetition”.
We investigated the repair behavior of the ten defective models that remained unresolved after six iterations. Among these models, five were categorized as “Counterexample,” four exhibited syntax errors, and one was identified as a “No instance”. Notably, the statuses of these ten specifications have remained largely unchanged since iterations 3 and 4. Particularly, the defective model arr1 consistently exhibited the same status from the first iteration onwards. This persistent behavior corroborates our preliminary analysis regarding the number of iterations, wherein no further progress was observed beyond the sixth iteration.
Figure 5: Repair progress in dual-agent Setting-6 (GPT-4-Turbo and Auto-Feedback) across iterations. Percentage numbers show transition volumes between statuses in subsequent iterations. Cumulative counts of fixed models follow each “Fixed” block.
Constraint-Related Issues. Our analysis of LLM-generated results reveals an important nuance in constraint-related issues: underconstrained and overconstrained specifications often coexist with various error manifestations. These fundamental constraint problems can simultaneously present as syntax errors, type errors, absence of valid instances, and counter examples. Rather than being distinct categories, constraint issues and error types frequently intersect. Addressing these interrelated challenges requires sophisticated specification engineering techniques that can balance constraint expressiveness with feasibility, ultimately leading to more reliable and robust LLM-based specification repairs.
Table 10: Frequency of overconstrained and underconstrained specifications among all unfixed issues across all LLM settings on ARepair and Alloy4Fun benchmarks.
ARepair Benchmark
Part A: Over-Constrained Issues
Model GPT-4-32k GPT-4-Turbo GPT-4o GPT-3.5-Turbo
1 2 3 4 5 6 7 8 9 10 11 12
addr 1 1 0 1 0 0 1 1 1 1 1 1
arr 1 1 1 0 1 0 0 0 0 1 0 0
balancedBST 1 1 0 1 1 1 1 1 1 2 2 1
bempl 0 0 0 0 0 0 0 1 1 0 0 0
cd 0 0 0 0 1 0 0 0 0 1 1 1
ctree 0 0 0 0 0 0 0 0 0 1 1 1
dll 0 0 0 1 1 0 0 0 0 2 0 0
farmer 0 0 0 0 0 0 0 0 0 0 0 0
fsm 0 0 0 1 1 0 0 0 0 1 1 1
grade 1 1 1 1 1 0 1 1 0 1 1 1
other 1 1 1 1 1 0 0 0 0 0 1 0
student 8 8 6 7 5 3 3 2 2 12 11 5
Summary 13 13 9 13 12 4 6 6 5 22 19 11
%ratio 56.5 59.1 56.3 61.9 63.2 40.0 42.9 54.5 50.0 64.7 59.4 55.0
Part B: Under-Constrained Issues
Model GPT-4-32k GPT-4-Turbo GPT-4o GPT-3.5-Turbo
1 2 3 4 5 6 7 8 9 10 11 12
addr 0 0 0 0 0 0 0 0 0 0 0 0
arr 1 1 1 0 1 0 0 0 0 1 0 0
balancedBST 1 1 0 1 1 1 1 1 1 1 1 1
bempl 1 1 1 0 0 0 1 0 0 1 1 1
cd 1 0 0 0 0 0 0 0 0 0 1 1
ctree 0 0 0 0 0 0 0 0 0 0 0 0
dll 0 0 0 1 1 0 0 0 0 1 1 1
farmer 0 0 0 0 0 0 0 0 0 0 0 0
fsm 1 1 0 1 1 1 1 1 0 1 1 1
grade 0 0 0 0 0 0 0 0 0 0 0 0
other 0 0 0 0 0 0 0 0 0 0 0 0
student 5 5 4 4 3 3 4 3 3 7 7 3
Summary 10 9 6 8 7 6 8 5 5 12 13 9
%ratio 43.5 40.9 37.5 38.1 36.8 60.0 57.1 45.5 50.0 35.3 40.6 45.0
Alloy4Fun Benchmark
Part A: Over-Constrained Issues
Model #issues Under Setting
GPT-4o GPT-3.5-Turbo
7 8 9 10 11 12
classroom (38) (39) (35) 612 610 593
cv (3) (1) (0) 60 65 41
graphs (23) (13) (23) 150 135 145
lts (13) (8) (9) 110 82 88
production (1) (0) (1) 22 18 16
trash (32) (33) (27) 120 120 114
Summary (110) (94) (95) 1074 1030 997
%ratio (60.8) (62.3) (61.0) 60.3 60.4 60.1
Part B: Under-Constrained Issues
Model #issues Under Setting
GPT-4o GPT-3.5-Turbo
7 8 9 10 11 12
classroom (22) (21) (19) 387 388 385
cv (1) (1) (1) 40 41 27
graphs (15) (7) (15) 110 97 103
lts (9) (4) (5) 71 52 55
production (0) (0) (1) 14 11 11
trash (24) (24) (20) 86 86 80
Summary (71) (57) (61) 708 675 661
%ratio (39.2) (37.7) (39.0) 39.7 39.6 39.9
Total ARepair Alloy4Fun
Issues 1 2 3 4 5 6 7 8 9 10 11 12 7 8 9 10 11 12
Over-constrained 13 13 9 13 12 4 6 6 5 22 19 11 (110) (94) (95) 1074 1030 997
Under-constrained 10 9 6 8 7 6 8 5 5 12 13 9 (71) (57) (61) 708 675 661
Our analysis proceeds in two principal steps to identify specification constraint defects across LLM-generated models. First, we detect overconstrained specifications by inspecting counterexamples and the presence of instances. If a counterexample is present and no instance is found—or if only a counterexample is detected, indicating overly restrictive constraints—we classify the model as overconstrained. This suggests that the model’s constraints are too strict to allow any valid instance. Second, for specifications not flagged as overconstrained, we identify underconstrained issues by comparing each model against our ground‐truth benchmarks. If there is a discrepancy in the constraints—such as a missing constraint or an implementation that fails to reflect the intended semantics—we label the model as underconstrained. This two‐stage procedure ensures that each model is categorized into exactly one defect class or deemed correct if neither condition applies.
Table 10 summarizes the distribution of overconstrained and underconstrained specifications among all unfixed issues across various LLM configurations on the ARepair and Alloy4Fun benchmarks, respectively. Notably, the results reveal that the constraint tendencies of each LLM can vary substantially depending on the benchmark context. On ARepair, GPT-4-32k and GPT-4-Turbo most frequently produce overconstrained specifications, with over 56% of unfixed issues falling into this category across most settings. GPT-4o shows mixed behavior, with some settings demonstrating a greater tendency toward underconstraining, particularly settings 7 and 9 (57.1% and 50.0% respectively).
In contrast, on the Alloy4Fun benchmark, GPT-4o consistently produces overconstrained specifications (60-62% of unfixed issues), while GPT-3.5-Turbo maintains a similar ratio of overconstrained issues (approximately 60%). This reversal in behavior between benchmarks is particularly notable, suggesting that its constraint tendencies are heavily influenced by the specific characteristics and complexities of the benchmark rather than being an intrinsic property of the model. These findings underscore a consistent trade-off between generating overly strict and overly permissive constraints, with distinct patterns emerging for each model and benchmark. Our systematic categorization method robustly captures these nuanced behaviors, providing insight into both model capabilities and the influence of benchmark design on LLM-driven specification repair.
@@
Analysis of failure characteristics during the repair process highlights common issues such as proposed repairs failing assertions and syntax errors, with observed transitions between failure categories over iterative prompting. Therefore, increasing the number of iterations does not necessarily result in a successful repair of the model.
5.4 Results for RQ4: Repair Costs
@@
We consider two primary cost factors to assess the financial implications of utilizing the APR pipeline with different LLMs: (i) time taken for bug resolution and (ii) monetary expenses associated with token consumption, based on OpenAI’s pricing as of March 2024.
@@
Time Analysis: We measure the time taken for bug resolution using the results presented in Table 5. Figure 6 illustrates the runtime analysis, contrasting fixed and unfixed scenarios. Across different settings, the median duration required for bug repair varies from 37.3 seconds in Setting-1 to 120.3 seconds in Setting-3. Notably, the maximum time spent on repairs reached 493.28 seconds in Setting-5. Conversely, the median duration associated with unfixed bugs amounted to 273.8 seconds and 392.2 seconds in Settings-1 and 4, respectively. Interestingly, all settings supporting the dual-agent setup (i.e., 3, 6, 9, and 12) consistently exhibit increased running time compared to the single-agent configuration, across all LLMs and under both fixed and unfixed bug scenarios.
@@
Figure 6: Distribution of running time across all settings for fixed bugs (left) and unfixed bugs (right).
@@
Monetary Costs: Figure 7 presents the monetary costs incurred under each setting for both fixed and unfixed models. Settings utilizing GPT-3.5-Turbo show the lowest costs for both fixed and unfixed models, attributed to the cheaper cost per token for GPT-3.5-Turbo compared to GPT-4 family models (see Table 3). Additionally, settings employing the dual-agent (3, 6, 9, and 12) exhibit marginally higher costs than their counterparts utilizing the same LLM under the single-agent setup.
@@
Figure 7: Monetary distribution across all settings for fixed bugs (left) and unfixed bugs (right).
The analysis of repair costs in RQ4 reveals notable variations in both execution time and monetary expenditure across different configurations. Among the models evaluated, GPT-3.5-Turbo consistently incurs the lowest cost for bug resolution. Within the GPT-4 family, GPT-4o demonstrates superior efficiency, achieving the lowest repair cost and runtime. Additionally, settings employing the dual-agent paradigm tend to exhibit slightly higher expenses compared to their single-agent counterparts.
6 Threats to Validity
@@
External Validity. The generalizability of our findings may be limited by the specific characteristics of the benchmarks used. To address this, we utilized diverse benchmark suites and evaluated the performance of the repair pipeline across various scenarios. Furthermore, we documented the experimental setup comprehensively to facilitate replication and external validation of our results. Finally, some repair instances in our benchmark have been fixed in prior work and may have been included in the training data of the LLMs. This raises the concern that the models might be memorizing existing repairs rather than generating novel solutions, which could limit their applicability to Alloy models they have not been trained on. As a result, this poses a threat to the generalizability of our findings, as the LLMs may perform well on known faults but struggle with previously unseen ones. Future work could further investigate this aspect, following the approach of Salerno et al. (2025), and introduce a new benchmark dataset—analogous to the ConDefects dataset Wu et al. (2024)—designed specifically to mitigate the issue of data leakage.
@@
7 Related Work
Several recent studies have explored the integration of Large Language Models (LLMs) into software engineering tasks, particularly in the realm of program repair.
@@ 7 Related Work
AlphaRepair Xia and Zhang (2022) leverages LLMs for APR in a zero-shot setting, but it requires removing the buggy line and replacing it with masked tokens. It then queries the LLM to fill-in the masked tokens with the correct tokens to generate patches. Xia et al. Xia and Zhang (2023b) improve APR performance by incorporating test feedback into prompts, while Kang et al. Kang et al. (2023) enabled LLMs to utilize a debugger for information gathering and patch generation. Additionally, Fuzz4All Xia et al. (2024) leverages LLMs as an input generation and mutation engine, employing an auto-prompting phase to generate concise input prompts. RepairAgent Bouzenia et al. (2024) employs LLMs, agents, and dynamic prompts for APR, albeit in the context of repairing Java applications. TestPilot Schäfer et al. (2024) uses LLMs to generate unit test cases for JavaScript, employing a few-shot learning approach to refine prompts with failed tests and error messages.
@@
In contrast to these efforts, our study focuses on using LLMs for automated specification repair, a less explored area in software engineering. We adopt an approach similar to Fuzz4All’s auto-prompting phase, specifically by incorporating LLM to construct the prompt Shin et al. (2020). Moreover, while existing work predominantly uses LLMs for program repair in imperative languages like Java and JavaScript, our study extends the application of LLMs to the domain of Alloy specifications, addressing a gap in the literature regarding specification repair techniques for declarative languages. Consequently, a direct quantitative comparison is largely infeasible, with the exception of the tool proposed by Hasan et al.Hasan et al. (2023), which utilizes LLMs to repair Alloy specifications. Nevertheless, Table 11 presents a comparative analysis of state-of-the-art LLM-based APR tools. This comparison highlights that our APR pipeline incorporates recent advances and techniques in LLMs that are also employed by contemporary repair tools.
Table 11: Comparing the feature that we considered in the APR pipeline with contemporary state-of-the-art LLM-APR tools, the column labeled “Autoprompt” denotes that the LLM generates the prompt.
APR Tools Agentic Tools Feedback Autoprompt Zero-shot Spec Repair # LLMs
AlphaRepair Xia and Zhang (2022) 1
ChatRepair  Xia and Zhang (2023a) 1
conversationalAPR Xia and Zhang (2023b) 10
Xia et al. Xia et al. (2023) 9
Hasan et al. Hasan et al. (2023) 1
RepairAgent Bouzenia et al. (2024) 1
Ours 4
8 Conclusion
@@
In this study, we explore the potential of pre-trained LLMs to facilitate the repair of Alloy specifications, taking into account recent advancements in LLMs, including the use of agents, feedback mechanisms, zero-shot learning capabilities, and auto-prompting techniques. The investigation reveals that employing a dual-agent repair pipeline enhances the repair process, albeit with a marginal increase in token consumption. The comparative evaluation highlights the superior performance of the GPT-4 model family over GPT-3.5-Turbo, underscoring their promising applicability. Overall, the findings of this research indicate a positive outlook for applying LLMs in automated program repair for declarative specifications, particularly through repair pipelines that integrate contemporary innovations in the field of LLMs.
8.1 Data Availability Statement
All data used to write this paper is open-source and publicly available at https://github.com/Mohannadcse/AlloySpecRepair.
Appendix: Additional Results
This section presents the results for the Alloy4Fun benchmark based on settings 7-12.
Figure 8: Venn diagrams showing the exclusive and overlapping successful repairs for the Alloy4Fun benchmark achieved by different repair methods. These illustrate the complementary and unique capabilities of the APR pipeline compared to state-of-the-art (SoTA) repair tools. Settings 7-9 are based on the sampled 357 models, while settings 10-12 are based on the entire set of models in the benchmark. Accordingly, the results are reported for the SoTA tools.
(a) Refer to caption
(b) Refer to caption
Figure 9: Alloy4Fun benchmark results: (a) Iteration count distribution for repairing specifications across various settings. (b) Incident count for initial repair attempts mirroring the buggy specification (lower values preferred).
Figure 10: Error types observed in failed repair iterations for the Alloy4Fun benchmark across settings 7-9.
Figure 11: Error types observed in failed repair iterations for the Alloy4Fun benchmark across settings 10-12.
Figure 12: Distribution of running time across settings 7-12 for the Alloy4Fun benchmark for fixed bugs (left) and unfixed bugs (right).
Figure 13: Monetary distribution across settings 7-12 for the Alloy4Fun benchmark for fixed bugs (left) and unfixed bugs (right).